# Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions

### v1

@@ II Results
Finally, we examine one important assumption of GHS, such that all households are willing to swap their home as long as the total commuting distance of the two swapping households decreases, even at the expense of an increase in the commuting distance of one of them. Here, we assume that the households with an increase of commuting distance have an $x\%$ willingness to swap home but $1-x\%$ to reject. As shown in Fig. S8 of the SI, there is no significant impact on $\Delta d$ as willingness $x$ changes.
@@
Reduction in Carbon Emissions. As a by-product, Greedy Home Swapping which reduces the average commuting distance and hence the traffic congestion, also reduces carbon emissions. In general, the longer the commuting distance, the longer the time vehicles stayed on the roads, the more congested were the roads, and the more emissions were produced  kissinger2019detailed ; xue2022impact ; muniz2018urban ; tian2024spatial ; bohm2022gross ; li2023assessing . To reveal the impact of home swapping on reducing carbon emissions, we examine the total carbon emissions (see Materials and Methods) in various cases as shown in Fig. 5A. As we can see, the total carbon emissions in RHS were much larger than the total empirical carbon emissions, and the total carbon emission in GHS was the least, such that the reduction in carbon emissions reaches $77.3\%$. We further compare road carbon emissions under RHS, real data, and GHS in rush hour in Fig. S9.
@@
For GHS with socio-demographic consideration, we see in Fig. 5A that carbon emissions are significantly higher, such that the total emissions are higher in c-GHS than those of p-GHS or a-GHS. The total emissions are similar in the case of p-GHS and a-GHS, and the total emissions after SDGHS under medium tolerance are much higher than that with a single constraint. As shown in the upper panel of Fig. 5C, after SDGHS with medium tolerance, specifically $T_{\rm center}$ = 0.5km, $T_{\rm price}$ = 15% and $T_{\rm amenity}$ = 0.3, the carbon emissions of some roads in the city where have decreased significantly compared to the emissions estimated by the original un-swapped data. When we consider only the constraints of $T_{\rm price}$ and $T_{\rm amenity}$, the lower panel of Fig. 5C shows much more roads with reduced carbon emissions. By comparing the insets of the upper and lower panels, we see that the reduction of emissions of the corresponding roads decrease significantly when the distance to the city center is considered in home-swapping in the upper panel, in line with the conclusion in Fig. 4F, implying that the distance to the city center plays a large role on GHS in reducing total emissions. Finally, we show the reduction in carbon emissions during the morning and the evening peak hour in each case in Fig. S10, such that similar results are obtained.
@@
Next, we analyze the dependence of the reduction in carbon emissions on the reduction in the average commuting distance, i.e. $\Delta d$, after GHS, as shown in Fig. 5B. As $\Delta d$ increases, the emission reduction continuously increases, with GHS reducing $30\%$ of the average commuting distance, leading to a $61\%$ reduction of carbon emissions. The nonlinear relation between the reduction in carbon emissions and the reduction in commuting distance has several origins. First, the changes in commuters’ departure time and route after home swapping change the congestion status of individual road segments. The reduction in congestion decreases the commuting time for a given commuting distance, thereby reduces carbon emissions. Additionally, as we have discussed before, when commuters change their residence locations, their type of travel also changes. When the commuting distance is less than 2.5 km, commuters may switch from driving to walking, leading to zero carbon emission for this group. We show the reduction of carbon emissions of individuals after GHS in Fig. S11, suggesting a potential trifurcation. Therefore, shortening average commuting distance does not only bring convenience to the residents’ daily lives but also hold a strong significance for protecting the urban environment and thus benefiting the health of the urban population.
@@
Exploiting GHS by Polycentric Urban Structure. As we know, polycentric city structure affects urban traffic nilforoshan2023human ; jun2020effects ; zuo2024interactions ; rong2022inspection ; louail2014mobile ; louf2013modeling ; pan2024whither ; sun2020effects ; zhao2024unravelling . The above analysis also suggests that the distance to the city center is the primary factor affecting the choice of the residence location of households and thus the reduction of commuting distance by home swapping. We thus study whether a polycentric urban structure could reduce significantly the average commuting distance. In this case, by shifting from a monocentric city model, where most activities are concentrated in a single region, to a polycentric model, with multiple centers of activity across the city. This approach could lead to a more balanced spatial distribution of living-working locations, allowing a more efficient use of transportation infrastructure, reducing travel times and carbon emissions, improving accessibility to amenities and services, and hence enhancing the quality of life for urban residents.
@@
Specifically, when SDGHS was conducted with medium tolerance, we found that $\Delta d$ increases with the number of city centers. To examine the impact of the distance between the city centers, we study the scenarios when the new centers are the most visited locations outward from the original city center at a minimum distance of 3km, 6km, 9km, and 12km, respectively, as shown in Fig. 6A. We found that $\Delta d$ increases with the distance between the city centers. For instance, in the case of a five-center city such that the four new centers are established at a distance of 12km outward from the original center, $\Delta d$ reaches $22.3\%$, suggesting that new urban centers should be established in peripheral areas to achieve a more effective reduction in commuting distance through home swapping. This suggests a strategic approach to urban planning whereas developing multiple activity centers with a distributed pattern and a larger distance between them can significantly improve commuting efficiency by home swapping, hence achieving a more efficient and energy-saving city. Similar result can be obtained in Shenzhen and Abidjan as shown in Fig. S19B and Fig. S23B respectively.
We further select bustling areas in the four peripheral counties of Shijiazhuang as new activity centers, to explore the impact of establishing multiple peripheral centers on the reduction in commuting distance. The results in Fig. 6B show that as the number of activity centers increases, the reduction of commuting distance increases; from a single center to dual centers, even without the corresponding supporting amenities, there is a significant reduction in commuting distance after SDGHS. From dual centers to multiple centers, the marginal increase in the reduction of commuting distance slows down. Other than the reduction in the average commuting distance, we show in Fig. 6C the congestion coefficient decreases with the number of activity centers, while carbon emissions are reduced after SDGHS under medium tolerance. This indicates that home swapping with the establishment of a polycentric city is a promising urban-planning solution in alleviating congestion and reducing carbon emissions.
@@ III Discussion
In this paper, we explored the benefit brought by an innovative approach to swap homes between households to reduce traffic congestion and carbon emissions in urban areas. By analyzing over 400,000 mobility trajectories in the Chinese city Shijiazhuang, we found that the overall city-wide commuting distance can be potentially reduced by over $50\%$. This finding is significant as it underscores the interplay between the distributions of citizens’ residence, workplace locations and urban traffic congestion. In addition, due to both a shorter commuting distance and reduced urban congestion, carbon emissions are also greatly reduced, which can be another important impact of home swapping. Our study is the first to verify the benefit brought by home-swapping using such a big real data.
Interestingly, to make home swapping more realistic, our analysis also revealed that when households considered the distance to the city center, housing price and amenity accessibility in swapping home, the reduction in commuting distance declines but still accounts to $13\%$, which is a substantial decrease. Moreover, among these factors, the distance from the city center emerged as the most influential factor for urban residents in choosing their residence location and whether to swap home. These results support the coordinated allocation of residence as an innovative strategy to substantially alleviate peak-hour traffic congestion and carbon emissions in densely populated metropolitans.
Though complete mandatory home swapping may not be readily feasible in reality, our findings are crucial in the broader context of urban planning and sustainable city development. By demonstrating the strong impact of residence allocation on traffic status, our work provides valuable insights and guidance for urban planners and policymakers in urban planning. It stresses the need for a holistic consideration on urban design, where residential planning is not only about housing city residents but also about integrating socio-economic factors including amenity accessibility, financial and environmental considerations. This study contributes significantly to the existing body of knowledge by providing empirical evidence on a unique, innovative yet determined home swapping approach on how urban structure can be coordinated and reformed, specifically towards a polycentric model to save energy, resources and time by significantly minimizing the fundamental source of traffic congestion. This is particularly relevant to tackle the challenges of both rapid urbanization and increasing urban population, in a board sense sheds light on sustainable city development.
While this study offers important and interesting insights, it has limitations. First, although the major analysis was based on a large dataset obtained in Shijiazhuang and similar results have been observed by using data in Shenzhen and Abidjan, analyses on more cities with different cultural, economic, or geographical characteristics can further extend its generalizability. Furthermore, the study primarily assumed residents commute by motor vehicles, not taking into account other modes of public transportation such as buses or metro. Future research could extend this work by examining a more diverse range of cities, incorporating multi-modal transportation data, and exploring the interplay between residence locations and different types of commuting behaviors. Additionally, investigating the psychological and social impacts of home swapping on individuals and communities would provide more understanding on this innovative approach, its feasibility and implications on urban planning and city development.

### v2

@@ II Results
Furthermore, we examine one important assumption of GHS, such that all households are willing to swap their homes as long as the total commuting distance of the two swapping households decreases, even at the expense of an increase in the commuting distance of one of them. Here, we assume that the households with an increase of commuting distance have an $w\%$ willingness to swap homes but $1-w\%$ to reject. As shown in Fig. S8A of the SI, there is no significant impact on $\Delta d$ as willingness $w\%$ changes.
In the above analysis, we focused on optimizing the commuting distance for households as a whole, without taking into account the individual preference within households. Here, if his/her commuting distance increases, we assume the commuter with an $s\%$ probability refuses to swap homes, which can be considered as individual selfishness in home swapping (see details in Section 3 in SI). Simulation results (Fig. S8B) suggest that the individual selfishness has little impact on $\Delta d$. However, in reality, residential decisions involve multiple factors beyond the distance to workplace. To take into account any potential requirements of the households, we randomly select a $p$ fraction of households who refuse to relocate regardless of the reduction in commuting distance, and swap homes only for the remaining 1p1-p of households. Simulation results of SDGHS show that even if $10\%$ to $90\%$ of households refuse to swap homes, a reduction of $12.7\%$ to $2.3\%$ of commuting distance is still achieved (see Fig. S9A).
To accommodate heterogeneous individual needs, we incorporate two principal constraints: access to educational facilities for children and the preservation of the shortest commuting distance within each household. First, considering the population of school-age children in Shijiazhuang, we randomly designate $18\%$ of households to refuse to swap homes, and in this case $\Delta d$ reaches $10.7\%$. To better align with real-world constraints, we analyze the spatial distribution of educational facilities (see Fig. S9B) and define households residing within the average street distance ($272$ m) of any school as “education-anchored”, restricting them from participating in home swapping. SDGHS in this case still achieves a $10.3\%$ reduction in commuting distance (see Fig. S9C).
Second, in another scenario, to account for the household member with the strongest need for a short commuting distance, for instance, caring for young children or the elderly, we refine home-swapping by defining a prioritized member in each household while maintaining system-wide benefits: home swapping are permitted if they (i) reduce the total commuting distance of both households to be swapped and (ii) maintain or reduce the commuting distance of the prioritized member in each household, defined as the one with the shortest commuting distance within the household. In this case, SDGHS achieves a reduction of $8.2\%$ of commuting distance. Under the combined constraints of educational resources and commuting priority, SDGHS still yielded an $8.1\%$ reduction in average commuting distance. All the above results demonstrate the robustness of the benefit of home swapping to real-world heterogeneity.
@@
Reduction in Carbon Emissions. As a by-product, Greedy Home Swapping which reduces the average commuting distance and hence the traffic congestion, also reduces carbon emissions. In general, the longer the commuting distance, the longer the time vehicles stayed on the roads, the more congested were the roads, and the more emissions were produced  kissinger2019detailed ; xue2022impact ; muniz2018urban ; tian2024spatial ; bohm2022gross ; li2023assessing . To reveal the impact of home swapping on reducing carbon emissions, we examine the total carbon emissions (see Materials and Methods) in various cases as shown in Fig. 5A. As we can see, the total carbon emissions in RHS were much larger than the total empirical carbon emissions, and the total carbon emissions in GHS was the least, such that the reduction in carbon emissions reaches $77.3\%$. We further compare road carbon emissions under RHS, real data, and GHS in rush hours in Fig. S10.
For GHS with socio-demographic consideration, we see in Fig. 5A that carbon emissions are significantly higher, such that the total emissions are higher in c-GHS than those of p-GHS or a-GHS. The total emissions are similar in the case of p-GHS and a-GHS, and the total emissions after SDGHS under medium tolerance are much higher than that with a single constraint. We show the reduction in carbon emissions during the morning and the evening peak hours in each case in Fig. S11, such that similar results are obtained. To more accurately reflect the consideration of households, we also analyze SDGHS with education-anchored households (SDGHS(e)), SDGHS with specific household members with the strongest need for the shortest commuting distance (SDGHS(s)), and SDGHS with the two combined considerations (SDGHS(es)). In these cases, the total emissions can be reduced by $34.4\%$, $27.6\%$ and $27.4\%$, respectively.
@@
As shown in the upper panel of Fig. 5C, after SDGHS with medium tolerance, specifically $T_{\rm center}=0.5km$, $T_{\rm price}=15\%$ and $T_{\rm amenity}=0.3$, the carbon emissions of some roads in the city where have decreased significantly compared to the emissions estimated by the original un-swapped data. When we consider only the constraints of $T_{\rm price}$ and $T_{\rm amenity}$, the lower panel of Fig. 5C shows much more roads with reduced carbon emissions. By comparing the insets of the upper and lower panels, we see that the reduction of emissions of the corresponding roads decrease significantly when the distance to the city center is considered in home-swapping in the upper panel, in line with the conclusion in Fig. 4F, implying that the distance to the city center plays a large role on home-swapping in reducing total emissions.
@@
Next, we analyze the dependence of the reduction in carbon emissions on the reduction in the average commuting distance, i.e. $\Delta d$, after GHS, as shown in Fig. 5B. As $\Delta d$ increases, the emission reduction continuously increases, with GHS reducing $30\%$ of the average commuting distance, leading to a $61\%$ reduction of carbon emissions. The nonlinear relation between the reduction in carbon emissions and the reduction in commuting distance has several origins. First, the changes in commuters’ departure time and route after home swapping change the congestion status of individual road segments. The reduction in congestion decreases the commuting time for a given commuting distance, thereby reduces carbon emissions. Additionally, as we have discussed before, when commuters change their residence locations, their type of travel also changes. When the commuting distance is less than 2.5 km, commuters may switch from motorized travel to non-motorized travel, leading to zero carbon emission. We show the carbon emissions of individuals after GHS in Fig. S12, suggesting a potential trifurcation. Therefore, shortening average commuting distance does not only bring convenience to the residents’ daily lives but also hold a strong significance for protecting the urban environment and thus benefiting the health of the urban population.
Based on the above analysis, we found that GHS and SDGHS can effectively reduce the total urban commuting distance, and alleviate the city-wide congestion and carbon emissions. However, considering the heterogeneity of individual, citizens may use public transport to commute a long distance instead of driving. In order to analyze the impact of multi-modal transportation, according to the commuting distance and average commuting speed based on the empirical data, we divide all commuters who use motorized travel into two modes, either by public buses or by private cars. As shown in Fig. S18A, the public transport system is established by the bus route data (see detail in Section S1 of SI). The results of GHS and SDGHS under multi-modal transportation were constructed in Fig. S18B and C, and it is indicated that $\Delta d$ and the reduction of carbon emissions are only slightly affected, meaning the inclusion of public transportation does not compromise the validity of the results of the proposed approach.
While our study primarily focuses on reducing commuting distance through home-swapping, we remark that additional non-commuting traffic may be induced owing to the reduced traffic congestion  duranton2011fundamental ; metz2021time ; loo2022spatio ; noland2006flow . We thus conduct analyses to demonstrate the robust benefit brought by home swapping even with non-commuting traffic induced: we add $0\%$-$25\%$ of trips exceeding $2.5$ km from the dataset during off-peak hours to the improved traffic patterns after GHS (see details in Sec. S8 of SI), such that the reduction in emissions decreases from $78.3\%$ to $71.7\%$ (see Fig. S17). Subsequent to the GHS, we also simulate SDGHS and SDGHS(es) incorporating $15\%$ of induced trips (i.e. roughly $63,000$ motorized trips). As denoted by iSDGHS and iSDGHS(es) in Fig. 5A, the reductions in carbon emissions are $39.8\%$ and $25.5\%$, respectively. This robust benefit from home swapping comes from the fact that the average distance of motorized commuters is $11.1$ km, higher than the average length of $7.6$ km of induced non-commuting trips. While the above simulation may not fully represent the amount and the pattern of induced trips, the strong periodicity of weekday travel patterns suggests induced trips in weekdays may be limited. These findings confirm the robustness of home swapping benefit and highlight an important direction for future research incorporating activity-based induced travel demand.
@@
Exploiting GHS by Polycentric Urban Structure. As we know, the polycentric city structure affects urban traffic congestion and carbon emissions nilforoshan2023human ; jun2020effects ; zuo2024interactions ; rong2022inspection ; louail2014mobile ; louf2013modeling ; pan2024whither ; sun2020effects ; zhao2024unravelling . The above analysis also suggests that the distance to the city center is the primary socio-demographic factor affecting the choice of the residence location of households and the reduction of commuting distance by home swapping. We thus study whether a polycentric urban structure could reduce significantly the average commuting distance. In this case, by shifting from a monocentric city model, where most activities are concentrated in a single region, to a polycentric model, with multiple centers of activity across the city. This approach could lead to a more balanced spatial distribution of living-working locations, allowing a more efficient use of transportation infrastructure, reducing travel time and carbon emissions, improving accessibility to amenities and services, and hence enhancing the quality of life for urban residents.
@@
Specifically, when SDGHS was conducted with medium tolerance, we found that $\Delta d$ increases with the number of city centers. To examine the impact of the distance between the city centers, we study the scenarios when the new centers are the most visited locations outward from the original city center at a minimum distance of 3 km, 6 km, 9 km, and 12 km, respectively, as shown in Fig. 6A. We found that $\Delta d$ increases with the distance between the city centers. For instance, in the case of a five-center city such that the four new centers are established at a distance of 12 km outward from the original center, $\Delta d$ reaches $22.3\%$, suggesting that new urban centers should be established in peripheral areas to achieve a more effective reduction in commuting distance through home swapping. This suggests a strategic approach to urban planning whereas developing multiple activity centers with a distributed pattern and a larger distance between them can significantly improve commuting efficiency by home swapping, hence achieving a more efficient and energy-saving city.
We further select bustling areas in the four peripheral counties of Shijiazhuang as new activity centers, to explore the impact of establishing multiple peripheral centers on the reduction in commuting distance. The results in Fig. 6B show that as the number of activity centers increases, the reduction of commuting distance increases; from a single center to dual centers, even without the corresponding supporting amenities, there is a significant reduction in commuting distance after SDGHS. From dual centers to multiple centers, the marginal increase in the reduction of commuting distance slows down. Other than the reduction in the average commuting distance, we show in Fig. 6C the average congestion coefficient decreases with the number of activity centers, while carbon emissions are reduced after SDGHS. This indicates that home swapping with the establishment of a polycentric city is a promising urban-planning solution in alleviating congestion and reducing carbon emissions.
@@
Universality of the benefit from home swapping. To verify the benefit of home swapping in cities other than Shijiazhuang, we studied an additional origin-destination dataset of commuters from another Chinese telecommunication operator, covering the morning traffic peak hour between 7-8 a.m. in 28 major Chinese cities (see Section S1 of SI for details). The absence of detailed individual trajectories in this data prevented the direct inference of household membership. Consequently, a simplified household matching mechanism was employed to construct households (see Section S11 of SI). First, we analyzed and found that in each of the 28 cities, the spatial distributions of commuters’ residence and workplace locations are similar, as observed in the case of Shijiazhuang, with an average KL divergence of $0.087$ (Fig. S20). This already indicates that home swapping is potentially beneficial in other major cities. As shown in Fig. 7A, we found that the average commuting distances decreases in these cities after GHS. Additionally, there is a slight positive correlation between $\Delta d$ and the city population size. Notably, Chongqing’s KL divergence is $0.067$, and $\Delta d$ after GHS is as high as $88\%$. The abnormally high $\Delta d$ might be caused by the mountainous geography of Chongqing which limit navigation. Likewise, this mountainous geography also makes the use of POI and housing price to describe regional (grid cells) characteristics invalid. To ensure the validity of the experiment, Chongqing is excluded from subsequent analysis.
By incorporating datasets from multiple sources, including amenities (POI data), housing prices, and commuting routes across the 28 cities, we further analyze the effectiveness of GHS with a single socio-demographic factor. In light of the characteristics of the datasets employed, adaptations are essential, involving the calibrated adjustment of key threshold parameters, i.e. the distance to city center with $T_{\rm center}=3km$, housing price with $T_{\rm price}=30\%$, and residential amenity accessibility with $T_{\rm amenity}=0.3$, to better align the model with empirical conditions. As shown in Fig. 7B, both p-GHS and a-GHS show similar $\Delta d$ and outperform c-GHS. The values of $\Delta d$ of each city, after c-GHS, a-GHS, and p-GHS, are shown in Fig. S21A.
By considering the joint effect of the three socio-demographic factors, Fig. 7C shows the values of $\Delta d$ after SDGHS with $T_{\rm center}=3km$, $T_{\rm price}=30\%$ and $T_{\rm amenity}=0.3$ for China’s top-10 congested cities conRep2023 . Although Chongqing ranks second among congested cities, due to Chongqing’s special mountainous geography as we mentioned above, Shijiazhuang replaces Chongqing in the following analysis. The results indicate that $\Delta d$ is typically smaller in cities with less severe urban congestion before home swapping, suggesting the significance of home swapping for megacities experiencing severe congestion. Furthermore, we examined the impact of polycentric urban structure on home swapping. Fig. 7E shows that polycentric cities achieved a greater reduction in average commuting distance after SDGHS than their monocentric counterparts (see details in Fig. S21B). To verify the positive correlation between $\Delta d$ and reduction in carbon emissions, we estimated the reduction in carbon emissions after GHS for the 10 cities (see Section S11 of the SI for details). As shown in Fig. 7D, a positive correlation exists between $\Delta d$ and carbon emissions reduction, consistent with the results from the study in Shijiazhuang using high-resolution trajectories (Fig. 5B). Collectively, this validation on multiple cities suggests that our findings do not depend on specific urban context but reflect fundamental principles of reducing commuting distance by home swapping. It is worth noting that the values of $\Delta d$ obtained by the origin-destination data after GHS are similar to the result obtained by the high-resolution trajectories in Shijiazhuang. However, the $\Delta d$ is significantly higher than that from Shijiazhuang when socio-demographic factors are considered. This discrepancy primarily originates from the coarse spatial granularity and the limited temporal resolution of the data, and an overly simplified household matching mechanism (see Section S11 of SI).
Data-driven model of residential choice. The above analysis shows that home swapping can effectively reduce the city-wide congestion and emissions. To explore the self-organizing mechanism of residential choice, and to analyze the feasibility of government coordination, we developed a data-driven model to describe residential choice (Fig. 8A) based on urban economics theory fujita1989urban and social physics yan2014universal ; barthelemy2019statistical . The model assumes that the attractiveness of location $i$ to household $j$ is dictated by three factors: (1) the housing affordability $E_{i}^{j}$, i.e. the alignment between the income level of household $j$ and the average housing price in the neighborhood of location $i$, (2) the locational utility $U_{i}$ of location $i$, such as amenities and infrastructure scaled by population density, and (3) the geometric mean commuting distance for household members $D_{i}^{j}$. Thus, the attractiveness $A_{i}^{j}$ of location $i$ for household $j$ can be expressed as
$A_{i}^{j}\propto\frac{E_{i}^{j}\times U_{i}}{D_{i}^{j}},$ (2)
where the expressions of $E_{i}^{j}$, $U_{i}$ and $D_{i}^{j}$ are shown in Materials and Methods. Based on this model and the empirical data of home-workplace location of urban households, we constructed a data-driven residential choice model for individual households. Simulation results show that if households choose their residential location according to this model, the total commuting distance at the city level is $1.1$ times that of the empirical data. Meanwhile, compared with the empirical morning peak average congestion coefficient ($0.75$), the model simulated average congestion coefficient is $0.78$.
To verify the model’s validity from the perspective of individual households, we compare the distribution of households’ average commuting distances obtained from the model with that of the empirical data. As shown in Fig. 8B, the blue dots represent model simulation results while the red dots show empirical data, demonstrating that the model’s statistical distribution aligns well with actual data. Furthermore, we compare the average commuting distance of individual household between model simulations ($\bar{d}_{\rm household}^{Model}$) and the empirical data ($\bar{d}_{\rm household}^{Real}$), as shown in the inset of Fig. 8B. The black dashed line indicates perfect equality between a household’s actual average commuting distance and model prediction, while the black solid lines represent cases where the difference between the empirical data and model simulation is less than 10 km. Statistics show that $78.2\%$ of individual households fall between these two black solid lines, indicating the model performs well at the individual household level. Therefore, this data-driven model outlines the principle and theory for analyzing residential location choice in the target city, which performs well at describing the self-organizing mechanism of residential choice.
To examine the compatibility of city-wide home-swapping within the urban economics framework, we extended the original model by incorporating government-coordinated interventions. Specifically, we introduced an economic subsidy policy where the government provides incentives for households to relocate to their guided optimal location of residence, which we call G-residence. The G-residence is defined as the residential location determined for each household through the GHS algorithm, exhibiting significant heterogeneity across different residential locations. Consequently, the relocation subsidies also vary substantially among households, as formalized in the following expression of policy incentive $\alpha_{i}^{j}$ for household $j$ to relocate to area $i$
$A_{i}^{j}\propto\alpha_{i}^{j}\times\frac{E_{i}^{j}\times U_{i}}{D_{i}^{j}},$ (3)
where the expressions of $\alpha_{i}^{j}$ is shown in the Materials and Methods, location $g$ is the government-recommended G-residence for household $j$ (as shown in Fig. 8A), and $\theta$ denotes the economic subsidy rate (ranging from $0\%$ to $100\%$) that household $j$ would receive when residing at location $g$. A higher $\theta$ value indicates stronger government intervention and greater government fiscal expenditure. To allow feasibility in this incentive provision scheme, we incorporate a distance-decay mechanism for the subsidy. If household $j$ resides in location $i$, they can also have a relative discount, which decays with the distance from location $i$ to G-residence (see details in Materials and Methods).
The reduction of average commuting distance ($\Delta d$) under different subsidy levels is shown in Fig. 8C. As the subsidy intensity $\theta$ increases, the city-wide commuting distance is further reduced. The results clearly demonstrate that in our urban economics-based data-driven residential choice model, with a policy-induced $\theta=10\%$, the total city-wide commuting distance achieves nearly $25\%$ reduction, while the morning peak average congestion coefficient decreases from $0.78$ to $0.55$, corresponding to a $55.46\%$ carbon emission reduction in morning peak hours.
In fact, many households might choose to reside in other location, far from the G-residence, even though the government provides a policy incentive. As shown in the inset of Fig. 8C, where the horizontal axis represents the policy input $\theta$ and the vertical axis shows the actual subsidy expenditure $\theta^{*}$ (i.e., the real financial input under the corresponding $\theta$ values), we observe a nonlinear relationship between $\theta^{*}$ and $\theta$. At $\theta=10\%$, the actual policy expenditure $\theta^{*}$ is only $6\%$. Fig. 8D visualize the city-wide road congestion condition during morning peak hours with $\theta=0\%$ and $\theta=10\%$, along with their respective average congestion coefficients. Comparing to Fig. 3E and Fig. 3F, the results clearly show that the introduction of subsidy for residential choice model achieves a reduction in commuting distance comparable to our proposed approach of city-wide home swapping without subsidy.
@@ III Discussion
In this paper, by analyzing over 400,000 mobility trajectories in the Chinese city of Shijiazhuang, we found the residential choice of households exhibits a self-organizing dynamics in terms of members’ commuting distances. Leveraging the observation of approximate distributions of residences and workplaces, we explored the potential benefit brought by a hypothetical home-swapping process between households to reduce traffic congestion and carbon emissions in urban areas. As proof of concept, the first proposed strategy of greedy home swapping (GHS) does not take socio-demographic factors and the needs of individual household members into account to reveal the upper bound of congestion and emissions reduction, and we found that the city-wide average commuting distance (measured by $\Delta d$) could be potentially reduced by over $50\%$. This substantial reduction underscores the critical relationship between the spatial distributions of citizens’ residences, workplace locations, and urban traffic congestion. To increase practical relevance, our analysis incorporated the socio-demographic factors, including the distance to the city center, housing price, and amenity accessibility in swapping homes, serving as proxies to characterize household-level heterogeneity. Compared with the results of GHS, the $\Delta d$ of socio-demographic constrained home swapping (SDGHS) declines but still accounts for $13.4\%$. Notably, among these factors, the distance to the city center emerged as the most influential factor for urban residents in choosing their residence location and deciding whether to swap homes. These results suggest that the coordinated allocation of residence with the development of a polycentric city is an innovative strategy to substantially alleviate peak-hour traffic congestion and carbon emissions in densely populated metropolitans. These findings were supported by a recent study with a small-scale household survey zhao2025housing , while the present study is the first to substantiate such benefits by guiding self-organizing residential choice using large-scale empirical trajectory data.
Nevertheless, both GHS and SDGHS swap homes by considering the household as a unit, neglecting the needs of individual household members. Thus, we further considered the dependence of home swapping on individual needs, especially the availability of schools for children and specific individual household members with the strongest need for a short commuting distance. By defining households within $272$ meters of schools as “education-anchored” families and restricting them from participating in home swapping, $\Delta d$ declines to $10.3\%$. On the other hand, prioritizing the need for the household members with the strongest desire for the shortest commute, $\Delta d$ reduces to $8.2\%$ after SDGHS. These results underscore the adaptability of our approach to heterogeneous residential preferences. While commuting distance may not always be the top priority, our framework can accommodate different considerations by households in their decision-making process on home swapping. Other than the reduction of traffic congestion, the combination of shorter commuting distance and less running time on roads (due to reduced urban congestion) leads to significant decreases in carbon emissions, highlighting another important benefit of home swapping. At the city level, without considering the household heterogeneity and individual needs, the reduction in carbon emissions reaches $77.3\%$ after GHS. Constrained by various factors, such as the socio-demographic factors and the needs of individual household members, carbon emissions can still be reduced by $27.4\%$-$34.4\%$.
While home swapping can greatly reduce the distance of commuting trips, the alleviation of traffic congestion may induce new non-commuting trips. By assuming there are $15\%$ new induced trips, and the three socio-demographic factors and the two individual needs are considered during the simulation of home swapping, the carbon emissions reduction reaches $25.5\%$ at the city level. The relation between the reduction in carbon emissions and the reduction in commuting distance reveals a nonlinear relationship, coming from the dual effect of the reduction of city-level commuting distances and street-level congestion. Development of the public transportation system is beneficial for reducing congestion and carbon emissions, and hence we also examined multi-modal transportation by integrating the city’s entire bus network and real-time schedules into our simulations. The results confirm that our proposed home-swapping approaches retain efficacy across diverse transportation modes. The robustness of our approach stems from the inherent flexibility of home-swapping in adapting to local constraints and household preferences, as well as to multi-modal transportation.
To verify the benefit from home swapping in cities other than Shijiazhuang, we used the commuter origin-destination dataset in 28 major Chinese cities to conduct home-swapping simulation. Similar findings on the benefit of home swapping are observed. The consistency between these large-scale simulations with coarse data and our high-precision single-city analysis suggests that our proposed method of home swapping has captured a fundamental urban management principle: coordinated residential choices can greatly maximize benefits at the individual and city levels. This foundational insight could effectively guide future theoretical research and inform policy development aimed at harmonizing micro-level behaviors with macro-scale urban objectives.
Finally, our residential choice model incorporates factors such as housing affordability, proximity to workplaces, and availability of amenities. This model, validated against empirical data, reproduces the observed distribution of households’ average commuting distances, thereby bridging our findings with utility-maximization principles. Further, we introduced policy interventions - such as a $10\%$ rent reduction - to explore how targeted incentives can facilitate or motivate the self-organized home swapping. Simulations reveal that even modest policy measures significantly enhance the redistribution of residential choice, accelerating reductions in city-wide congestion and emissions. Through this analysis, we demonstrate that home swapping with incentives of subsidies possesses substantial theoretical feasibility within urban economics theory, which clarifies how our approach complements - rather than contradicts - urban economics theory, offering a practical pathway to integrate sustainability goals into decentralized residential choice dynamics.
We acknowledge the practical challenges of home-swapping on relocation costs and behavioral inertia. However, by demonstrating that voluntary swaps driven by shared economic and environmental incentives can reshape commuting patterns, our findings provide a proof of concept for market-compatible policies (e.g., congestion pricing rebates, green housing tax credits, etc.) that indirectly promote similar self-organization. Our work emphasizes the need for a holistic consideration of urban design, where residential planning is not only about housing city residents but also about integrating socio-economic factors including, amenity accessibility, financial and environmental considerations. We also remark that the reallocation of residences is a short-term loss, while the reduction in commuting distance is a long-term gain, and hence the gain must outweigh the loss in the long run. In addition, while city-wide home swapping would be costly, our work identifies guiding principles for incremental, incentive-driven policies that align individual and systemic goals. More importantly, our model simulations reveal that even with the modest interventions - such as a $6\%$ effective fiscal outlay via targeted rent subsidies - one could achieve a $25\%$ total commuting distance reduction and a $55\%$ emission reduction by nudging households toward mutually beneficial swaps. These nonlinear returns suggest that partial, phased adoption of home-swapping (e.g., pilot districts, priority corridors, etc.) could yield disproportionate sustainability gains without requiring abrupt, city-wide reorganization. Thus, this study contributes significantly to the existing body of knowledge by providing empirical evidence on a unique, innovative, yet determined home-swapping approach on how urban structure can be coordinated and reformed, specifically towards a polycentric model to save energy, resources, and time by significantly minimizing the fundamental source of traffic congestion. This is particularly relevant to tackling the challenges of both rapid urbanization and increasing urban population and, in a broad sense, sheds light on sustainable city development.
While this study offers important and interesting insights, it has limitations. First, although the major analysis was based on a dataset of high-resolution trajectories obtained in Shijiazhuang and similar results have been observed by using origin-destination data covering 28 major Chinese cities, examining more cities with different cultures and economic backgrounds would improve its generalizability. Second, in cities with distinctive geographic complexity, such as Chongqing, conventional evaluation methods become inadequate, potentially compromising the accuracy of an accurate assessment of the benefit brought by home-swapping. Furthermore, some other special personal preferences, on top of school districts and caregiving needs which we have already included in our study, may affect the choice of residential locations. Some of these factors are challenging to quantify or to obtain due to privacy and the absence of standardized city-level metrics, limiting us from taking these factors into account in our study. Additionally, the regional urbanization trajectories, governance structures, local administrative style, and cultural norms (e.g., homeownership preferences) of the target city could also influence implementation. Finally, the analysis acknowledges the potential for a rebound effect, where reduced peak-hour commuting traffic may induce additional non-commuting trips, which could offset the benefits of home-swapping. This underscores the critical role of complementary government policies to manage induced travel demand in ensuring the feasibility and effectiveness of home-swapping.